# Test-Time Adaptation with State-Space Models

### v1

@@ 5.1 Natural Temporal Distribution Drifts
Table 1: Accuracy on Wild-Time benchmarks averaged over three random training seeds
Methods Yearbook FMoW-Time
Source Model 81.30 $\pm$ 4.18 68.94 $\pm$ 0.20
Batch Norm (BN) 84.54 $\pm$ 2.10 10.14 $\pm$ 0.04
TENT 84.53 $\pm$ 2.11 10.21 $\pm$ 0.01
CoTTA 84.35 $\pm$ 2.13 10.19 $\pm$ 0.04
STAD-vMF 85.50 $\pm$ 1.30 86.25 $\pm$ 1.18
STAD-vMF + BN 86.20 $\pm$ 1.23 9.26 $\pm$ 1.97
STAD-Gaussian 86.22 $\pm$ 0.84
STAD-Gaussian + BN 86.56 $\pm$ 1.08
We start by evaluating the adaptation abilities of STAD to natural temporal drifts on two image classification datasets of the Wild-Time benchmark [40]. While only a binary problem, the Yearbook task is difficult as the images are low-resolution (32-dimensional feature space). FMoW-Time presents an even more challenging setting: 62 classes, high-resolution images, 1024-dimensional feature space. For Yearbook, we report both the vMF and Gaussian STAD variants. For FMoW-Time, the high-dimensional representation space makes the Gaussian model too computationally costly so we evaluate just the vMF version. For Yearbook we report results for a batch size of 2048 comprising all images of a year in one batch; for FMoW, we re-use the training batch size of 64.
STAD reliably adapts to natural shifts over time
Table 1 shows overall accuracy, averaged over all time steps and three random seeds. Our methods best adapt to the natural temporal shift present in both datasets, improving upon the source model in both cases. Strikingly, traditional CTTA methods collapse on FMoW-Time, falling well below the source accuracy by over 58 percentage points. On the other hand, vMF adaptation gains over 17 percentage points on the source model. This is an example of when adapting a powerful feature extractor such as the DenseNet [18] pretrained on ImageNet [11] can lead to catastrophic forgetting. On Yearbook, both Gaussian and vMF outperform baselines with the fully parameterized Gaussian model expectedly modeling the distribution shift better than the light-weight vMF model. We also test our last layer adaptation in combination with BN which modifies the feature extractor. This additionally, yields a marginal performance increase of < 1% point on Yearbook. Figure 3 (right) displays adaptation performance over different timestamps.
Figure 3: Adaptation performance on Wild-Time benchmarks: First column: Accuracy for different batch sizes. STAD is more robust to small batch sizes compared to baselines. Right column: Adaptation accuracy over different test time points. STAD reliably adapts to the distribution shifts. Baselines perform similarly, which is depicted by overlaying accuracy trajectories.
STAD excels on small batch sizes
Previous work [24, 41, 28] has observed that test-time adaptation methods are quite sensitive to the size of the test batch. To assess the impact of batch sizes, we also perform adaptation on Yearbook and FMoW using smaller batch size values. Figure 3 shows accuracy as a function of test batch size. Leveraging Bayesian principles, STAD is extremely robust to the test batch size, improving upon the source model across almost all batch sizes. In contrast, BN, TENT and CoTTA harm the source model when batch sizes are smaller than 512 on Yearbook and fail to improve entirely for FMoW. Adapting in small-sample environments is particularly crucial in continual domain adaptation when the task requires that predictions be made quickly. We also find that STAD is more robust to class imbalance (Section D.1) and study runtime in Section D.2.
5.2 Ablation Study: How important is modeling the dynamics?
Table 2: Difference in accuracy between dynamic and static versions of STAD (i.e.when removing the transition model). Performance drops substantially when dynamics are removed from the model.
Variant Yearbook FMoW CIFAR-10-C
STAD-vMF w/o dynamics –24.47 –17.38 –3.41
STAD-Gaussian w/o dynamics –28.43
We next investigate the importance of STAD’s temporal component. STAD is proposed with the assumption that adapting the class prototypes based on those of the previous time step facilitates both rapid and reliable adaptation. However, one could also consider a static version of STAD that does not have a transition model (Equation 1). Rather, the class prototypes are computed as a standard mixture model (Equation 2) and without considering previously inferred prototypes. Table 2 presents the accuracy differences between the static and dynamic versions of STAD in percentage points. Removing STAD’s transition model results in a substantial performance drop of up to 28 percentage points. This supports our assumption that state-space models are well-suited for the task of continual adaptation.
5.3 Limitation: Representations are a bottleneck
We next turn to the CIFAR-10-C dataset, which is the most commonly used benchmark in the CTTA literature. Table 3 displays adaptation accuracy. STAD improves upon the source model for all severity levels, but does not outperform TENT or CoTTA. We suspect this reversal in performance (i.e. STAD performing better than baselines on natural but worse on synthetic shifts) is related to the quality of the source model’s representations. Recall that, unlike TENT, CoTTA or BN, STAD adapts only the model’s last layer. While this has benefits for STAD’s wide applicability, it is also a limitation in that STAD can only be as good as the last-layer representations allow.
Figure 4: Analysis of cluster structure on CIFAR-10-C: Left: Dispersion (angular degrees) decreases with corruption severity, causing the ground truth cluster centers to become more similar. Middle: The angular distance to the ground truth cluster centers is smaller for STAD compared to the source model. Right: Dispersion of inferred prototypes can predict when adaptation with STAD is discouraged.
Table 3: Accuracy on CIFAR-10-C for different severity levels averaged over all corruption types
@@ 5.3 Limitation: Representations are a bottleneck
STAD-vMF 88.21 83.68 78.42 72.19 62.44 76.99
STAD-vMF + BN 90.16 88.11 86.24 83.17 79.33 85.40
To investigate representation quality as the cause for this performance drop, we employ the dispersion metric [27] developed for hyperspherical features: $\mathrm{dis}({\mathbf{W}}_{t})=\frac{1}{K}\sum_{k=1}^{K}\frac{1}{K-1}\sum_{l\neq k}^{K}{\mathbf{w}}^{T}_{t,k}{\mathbf{w}}_{t,l}$, where ${\mathbf{w}}_{t,k}$ denotes the class prototypes. Dispersion measures how far apart prototypes of different classes are. High dispersion values (in angular degrees) indicate prototypes point in different directions and thus are desirable. Figure 4 (left) shows dispersion on CIFAR-10-C. As corruption severity increases (x-axis), the dispersion of STAD’s representations (blue line) decreases. However, this is not due to any mis-estimation problem with STAD as the ground truth representations (yellow line)—computed using the true test labels—also decrease in quality. In Figure 4 (middle), we confirm that STAD is appropriately adapting by measuring the angular distance to the ground-truth representations. STAD (blue line) is substantially closer to the ground truth than the source model (gray line) is, showing that STAD well tracks the class prototypes as they evolve.
Dispersion measures the quality of representation, making it a potential diagnostic tool. If dispersion is high, STAD is all that is needed for good performance. If dispersion is low, a full-model adaptation strategy such as BN is required. In Figure 4 (right), we plot accuracy vs dispersion, showing that they positively correlate. This enables real-time signaling of insufficient representation quality, alerting the user to potential adaptation risks without the need for labels. Recognizing when adaptation fails is of critical importance in practice. Recent work has shown that most CTTA methods eventually collapse [31, 32], but understanding why and when this phase transition occurs is an open problem.
@@ 6 Conclusion
We have presented a novel test-time adaptation strategy, State-space Test-time ADaptation (STAD), based on a probabilistic state-space model. Both our Gaussian and vMF variants of STAD track how the last-layer evolves under distribution shift, allowing a deployed model to perform unsupervised adaptation to the shift. Our framework outperforms state-of-the-art competitors such as TENT and CoTTA on Wild-Time benchmarks. Yet we also identify points of improvement, as we found that random corruptions applied to CIFAR-10 degraded the representations to a degree that our method could be competitive with but not outperform these baselines adapting the model backbone. For future work, we will study diagnostics that reliably identify when the last-layer representations are suitable, more intensive adaptation is needed or if adaptation is possible at all. We will also investigate non-linear models of shift dynamics.
Acknowledgments and Disclosure of Funding
@@
  • Appendix A provides details on the Gaussian formulation of STAD (STAD-Gaussian) stating the employed transition and emission model.
  • Appendix B details the inference for the von Mises-Fisher formulation (STAD-vMF) listing the update equations for the variational EM step.
  • Appendix C contains various implementation details regarding our experiments on Yearbook and FMoW (Section C.1) and CIFAR-10-C (Section C.2).
  • Appendix D shows additional experimental results on adaptation under class imbalance (Section D.1) and presents runtime comparisons (Section D.2).
Appendix A STAD-Gaussian
We use a linear Gaussian transition model to describe the weight evolution over time: For each class $k$, the weight vector evolves according to a linear drift parameterized by a class-specific transition matrix ${\bm{A}}_{k}\in{\mathbb{R}}^{D\times D}$. This allows each class to have independent dynamics. The transition noise follows a multivariate Gaussian distribution with zero mean and global covariance $\bm{\Sigma}^{\mathrm{trans}}\in{\mathbb{R}}^{D\times D}$. The transition noise covariance matrix is a shared parameter across classes and time points to prevent overfitting and keep parameter size at bay. Equation 12 states the Gaussian transition density.
@@ Appendix A STAD-Gaussian
Equation 13 gives the emission model of the observed features ${\mathbf{H}}_{t}$ at time $t$. As in Equation 2, the features at a given time $t$ are generated by a mixture distribution with mixing coefficient $\pi_{t,k}$. The emission density of each of the $K$ component is a multivariate normal with the weight vector of class $k$ at time $t$ as mean and $\bm{\Sigma}^{\mathrm{ems}}\in{\mathbb{R}}^{D\times D}$ as class-independent covariance matrix. The resulting model can be seen as a mixture of $K$ Kalman filters. Variants of it has found application in applied statistics [6].
@@
We use the EM objective of Equation 5 to maximize for the model parameters $\phi=\{\{{\bm{A}}_{k},\{\pi_{t,k}\}_{t=1}^{T}\}_{k=1}^{K},\bm{\Sigma}^{\mathrm{trans}},\bm{\Sigma}^{\mathrm{ems}}\}$. Thanks to the linearity and Gaussian assumptions, the posterior expectation $\mathbb{E}_{p({\mathbf{W}},{\mathbf{C}}|{\mathbf{H}})}[\cdot]$ in Equation 5 can be computed analytically using the well known Kalman filter predict, update and smoothing equations [6, 5].
@@
Appendix B Inference for STAD-vMF
@@
Using the von Mises-Fisher distribution as hyperspherical transition (Equation 8) and emission model (Equation 9), the log of the complete-data likelihood in Equation 3 becomes
@@
where $\kappa_{0,k}$ and $\bm{\mu}_{0,k}$ denote the parameters of the first time step. In practise, we set $\bm{\mu}_{0,k}$ to the source weights and $\kappa_{0,k}=100$ (see Appendix C).
@@
As described in Section 3.3, we approximate the posterior $p({\mathbf{W}}_{1:T},{\mathbf{C}}_{1:T}|{\mathbf{H}}_{1:T})$ with a variational distribution $q({\mathbf{W}}_{1:T},{\mathbf{C}}_{1:T})$ assuming the factorised form
@@
Taking the expectation of the complete-data log likelihood (Equation 14) with respect to the variational distribution (Equation 20) gives
@@
Maximizing objective (Equation 22) with respect to the model parameters $\phi=\{\kappa^{\mathrm{trans}},\kappa^{\mathrm{ems}},\{\pi_{t,k}\}_{t=1}^{T}\}_{k=1}^{K}\}$ gives
@@
Here we made use of the approximation from Banerjee et al. [3] to compute an estimate for $\kappa$,
@@ Appendix C Experimental Details
C.1 Yearbook and FMoW
We follow the Eval-Fix protocol of Wild-Time to assess adaptation performance on Yearbook and FMoW. In Eval-Fix, source models are trained on a fixed time period and evaluated on unseen, future time points. We use the models trained by Yao et al. [40] on the training time period and evaluate them on the test period. Yao et al. [40] uses a range of different training procedures. We use the checkpoints for plain empirical risk minimization. Three different random training seeds are provided by Yao et al. [40]. All adaptation methods are continuously running without resets. Each method uses one optimization step (via entropy minimization for TENT and CoTTA and Expectation-Maximization for STAD).
By the nature of test-time adaptation, choosing hyperparameters is difficult since one cannot assume access to a validation set of the test distribution in practise. To ensure we report the optimal performance of baseline models, we conduct a grid search on the test set for relevant hyperparameters and report the performance of the best setting. The hyperparameters tested as well as other configurations are listed next.

### v2

@@ 5.2 Beyond Temporal Shifts: Reproduction Datasets and Synthetic Corruptions
Table 3: Accuracy on reproduction datasets and label shifts, averaged over three random data seeds.
CIFAR-10.1 ImageNetV2
Method covariate shift + label shift covariate shift + label shift
Source model 88.25 63.18
adapt feature extractor
BN 86.45 $\pm$ 0.28 23.83 $\pm$ 0.31 62.69 $\pm$ 0.15 43.20 $\pm$ 0.28
TENT 86.75 $\pm$ 0.35 23.87 $\pm$ 0.06 63.00 $\pm$ 0.16 43.20 $\pm$ 0.28
CoTTA 86.75 $\pm$ 0.17 22.37 $\pm$ 0.25 61.66 $\pm$ 0.29 43.73 $\pm$ 0.33
SHOT 86.50 $\pm$ 0.23 23.83 $\pm$ 0.31 62.97 $\pm$ 0.22 43.10 $\pm$ 0.34
SAR 86.45 $\pm$ 0.28 23.82 $\pm$ 0.33 62.99 $\pm$ 0.10 43.19 $\pm$ 0.25
adapt classifier
LAME 88.20 $\pm$ 0.09 92.42 $\pm$ 0.28 63.15 $\pm$ 0.10 80.47 $\pm$ 0.32
T3A 88.28 $\pm$ 0.06 89.00 $\pm$ 0.66 62.86 $\pm$ 0.04 63.47 $\pm$ 0.09
STAD-vMF (ours) 88.42 $\pm$ 0.10 92.23 $\pm$ 0.70 62.39 $\pm$ 0.05 81.46 $\pm$ 0.24
Reproduction Datasets
We evaluate our method on reproduction datasets (CIFAR-10.1 and ImageNetV2), which have recently gained attention as benchmarks for more realistic and challenging distribution shifts (Zhao et al., 2023b). We use a batch size of 100 for CIFAR-10.1 and 64 for ImageNetV2. Tab. 3 confirms the difficulty of adapting to more natural distribution shifts. For CIFAR-10.1, only T3A and STAD outperform the source model for both with and without label shift, with STAD adapting best. For ImageNetV2, none of the methods improve upon the source model when the label distribution is uniform. We again observe that classifier-adaptation methods handle label shifts better by a significant margin.
Table 4: Accuracy on CIFAR-10-C across different levels of synthetic corruptions
@@
SHOT 90.31 88.66 87.31 85.02 82.13 86.69
SAR 90.16 88.09 86.26 83.32 79.48 85.46
adapt classifier
LAME 86.94 81.39 74.93 67.69 56.47 73.48
T3A 87.83 82.75 76.77 69.43 57.90 74.94
STAD-vMF (ours) 88.21 83.68 78.42 72.19 62.44 76.99
Synthetic Corruptions
Lastly, we test our method on gradually increasing noise corruptions of CIFAR-10-C, a standard TTA benchmark. We use a batch size of 100. Tab. 4 shows the accuracy averaged across all corruption types. We make three key observations. First, performance gains are much higher than on previous datasets indicating the challenge posed by non-synthetic shifts. Second, as expected, methods adapting the backbone model are more performative on input-level noise, since such shifts primarily affect earlier layers (Tang et al., 2023; Lee et al., 2023). Lastly, STAD is consistently the best method amongst those adapting only the last linear layer.
5.3 Analysis of Tracking Abilities
Lastly, we seek to further understand the reasons behind STAD’s strong performance. At its core, STAD operates through a mechanism of dynamic clustering. We next inspect the importance of STAD’s dynamics component and assess the fidelity of its clustering.
Clusters are reliable.
Figure 6: Cluster fidelity on CIFAR-10-C
We evaluate how well STAD’s inferred cluster centers align with the ground truth cluster centers (computed using labels). We chose the progressively increasing distribution shift of CIFAR-10-C as this dataset represents a bigger challenge for STAD. Fig. 6 (left) shows distance (in angular degrees) to the ground truth cluster centers for both the source model and STAD. STAD (blue line) adapts effectively, significantly reducing the angular distance to the ground truth cluster centers. For the source model, the progressive distribution shift causes the ground truth cluster centers to drift increasingly further from the source prototypes (grey line). Additionally, by computing dispersion (Ming et al., 2023) (Fig. 6, middle), which measures the spread of the prototypes (in angular degrees), we find that STAD mirrors the ground truth trend (yellow line) of clusters becoming closer together. This is a promising insight, as it suggests that STAD’s cluster dispersion could potentially serve as an unsupervised metric to proactively flag when clusters start overlapping and estimate adaptation accuracy. In Fig. 6 (right), we plot accuracy vs dispersion of STAD’s prototypes for different corruptions and severity levels, confirming that they positively correlate.
Dynamics are crucial.
Table 5: Accuracy of dynamic and static versions of STAD (i.e. when removing the transition model)
Variant Yearbook FMoW CIFAR-10-C
STAD-vMF with dynamics 85.50 $\pm$ 1.30 86.25 $\pm$ 1.18 76.99
STAD-vMF w/o dynamics 61.03 $\pm$ 2.92 68.87 $\pm$ 0.28 73.57
Delta –24.47 –17.38 –3.41
STAD-Gauss with dynamics 86.22 $\pm$ 0.84
STAD-Gauss w/o dynamics 57.79 $\pm$ 2.14
Delta –28.43
STAD is proposed with the assumption that adapting the class prototypes based on those of the previous time step facilitates both rapid and reliable adaptation. However, one could also consider a static version of STAD that does not have a transition model (Eqn. 1). Rather, the class prototypes are computed as a standard mixture model (Eqn. 2) and without considering previously inferred prototypes. Tab. 5 presents the accuracy differences between the static and dynamic versions of STAD in percentage points. Removing STAD’s transition model results in a substantial performance drop of up to 28 points. This supports our assumption that state-space models are well-suited for the task of TempTTA.
@@ 6 Conclusion
We have presented State-space Test-time ADaptation (STAD), a novel test-time adaptation strategy based on probabilistic state-space models that addresses the challenges of temporal distribution shifts. Our Gaussian and vMF variants of STAD effectively track the evolution of the last layer under distribution shifts, enabling unsupervised adaptation in deployed models. Our extensive experiments highlight the significant challenges that temporal distribution shifts present for existing TTA methods. Future work on TempTTA could explore incorporating temporal information, such as timestamps, to better model the passage of time and extend the approach to irregularly sampled settings.
Acknowledgments
@@
  • App. A provides details on the Gaussian formulation of STAD (STAD-Gauss) stating the employed transition and emission model.
  • App. B details the von Mises-Fisher formulation (STAD-vMF). In Sec. B.1, we provide the inference scheme including all update equations for the variational EM step. In Sec. B.2, we provide the algorithmic overview of STAD-vMF.
  • App. C contains various implementation details.
  • App. D shows additional experimental results on single sample adaptation (Sec. D.1), STAD in combination with BN (Sec. D.2), visualizations of representation space (Sec. D.3) runtime comparisons (Sec. D.4) and sensitivity to hyperparameters (Sec. D.5).
Appendix A Details on STAD-Gaussian
A.1 Model Formulation
We use a linear Gaussian transition model to describe the weight evolution over time: For each class $k$, the weight vector evolves according to a linear drift parameterized by a class-specific transition matrix ${\bm{A}}_{k}\in{\mathbb{R}}^{D\times D}$. This allows each class to have independent dynamics. The transition noise follows a multivariate Gaussian distribution with zero mean and global covariance $\bm{\Sigma}^{\mathrm{trans}}\in{\mathbb{R}}^{D\times D}$. The transition noise covariance matrix is a shared parameter across classes and time points to prevent overfitting and keep parameter size at bay. Eqn. 10 states the Gaussian transition density.
@@ A.1 Model Formulation
Eqn. 11 gives the emission model of the observed features ${\mathbf{H}}_{t}$ at time $t$. As in Eqn. 2, the features at a given time $t$ are generated by a mixture distribution with mixing coefficient $\pi_{t,k}$. The emission density of each of the $K$ component is a multivariate normal with the weight vector of class $k$ at time $t$ as mean and $\bm{\Sigma}^{\mathrm{ems}}\in{\mathbb{R}}^{D\times D}$ as class-independent covariance matrix. The resulting model can be seen as a mixture of $K$ Kalman filters. Variants of it has found application in applied statistics (Calabrese & Paninski, 2011).
@@
We use the EM objective of Eqn. 5 to maximize for the model parameters $\phi=\{\{{\bm{A}}_{k},\{\pi_{t,k}\}_{t=1}^{T}\}_{k=1}^{K},\bm{\Sigma}^{\mathrm{trans}},\bm{\Sigma}^{\mathrm{ems}}\}$. Thanks to the linearity and Gaussian assumptions, the posterior expectation $\mathbb{E}_{p({\mathbf{W}},{\mathbf{C}}|{\mathbf{H}})}[\cdot]$ in Eqn. 5 can be computed analytically using the well known Kalman filter predict, update and smoothing equations (Calabrese & Paninski, 2011; Bishop & Nasrabadi, 2006).
@@
Appendix B Details on STAD-vMF
B.1 Inference
@@
Using the von Mises-Fisher distribution as hyperspherical transition and emission model, the log of the complete-data likelihood in Eqn. 3 becomes
@@
where $\kappa_{0,k}$ and $\bm{\mu}_{0,k}$ denote the parameters of the first time step. In practise, we set $\bm{\mu}_{0,k}$ to the source weights and $\kappa_{0,k}=100$ (see App. C).
@@
As described in Sec. 3.3, we approximate the posterior $p({\mathbf{W}}_{1:T},{\mathbf{C}}_{1:T}|{\mathbf{H}}_{1:T})$ with a variational distribution $q({\mathbf{W}}_{1:T},{\mathbf{C}}_{1:T})$ assuming the factorised form
@@
Taking the expectation of the complete-data log likelihood (Eqn. 12) with respect to the variational distribution (Eqn. 18) gives
@@
Maximizing objective (Eqn. 20) with respect to the model parameters $\phi=\{\kappa^{\mathrm{trans}},\kappa^{\mathrm{ems}},\{\pi_{t,k}\}_{t=1}^{T}\}_{k=1}^{K}\}$ gives
@@
Here we made use of the approximation from Banerjee et al. (2005) to compute an estimate for $\kappa$,
@@
B.2 Algorithmic Overview
Algorithm 1 STAD-vMF
1:Input: source model $f_{\theta}$, test batches $\{{\mathbf{X}}_{t}\}_{t=1}^{T}$, sliding window size $s$
2:Initialize: mixing coefficients $\pi_{t,k}\leftarrow\frac{1}{K}\forall t,k$, weights ${\mathbf{W}}_{t}\leftarrow{\mathbf{W}}_{0}\forall t$, transition concentration $\kappa^{trans}$ and emission concentration $\kappa^{ems}$
3:for $t\in\mathcal{T}$ do
4:     Define sliding window $S_{t}=\{\tau\mid\max(1,t-s)\leq\tau\leq t\}$
5:     Get representations ${\mathbf{H}}_{t}\leftarrow f_{\theta}^{L-1}({\mathbf{X}}_{t})$
6:     E-step:
7:     for $\tau\in S_{t}$ do
8:         for $k=1,\dots,K$ do
9:              for $n=1,\dots,N_{t}$ do
10:                  Compute $\mathbb{E}[c_{n,\tau,k}]$ by Eqn. 27 $\triangleright$ Cluster assignments
11:              end for
12:              Compute $\mathbb{E}[{\mathbf{w}}_{\tau,k}]$ by Eqn. 28 $\triangleright$ Prototype dynamics
13:         end for
14:     end for
15:     M-Step:
16:     for $\tau\in S_{t}$ do
17:         Compute $\pi_{\tau}$ by Eqn. 31 $\triangleright$ Label distribution
18:     end for
19:     Compute $\kappa^{\mathrm{ems}}$ by Eqn. 30 (optional)
20:     Compute $\kappa^{\mathrm{trans}}$ by Eqn. 29 (optional)
21:     Predict:
22:     Get predictions $\{y_{t,n}\}_{n=1}^{N_{t}}$ by Eqn. 6
23:end for
24:return Predictions $\{y_{t,n}\}_{n=1}^{N_{t}}$
@@ Appendix C Experimental Details
C.1 Source Architectures
  • CNN: We employ the four-block convolutional neural network trained by Yao et al. (2022) to perform the binary gender prediction on the yearbook dataset. Presented results are averages over three different seeds trained with empirical risk minimization. The dimension of the latent representation space is 32.
  • WideResNet: For the CIFAR-10 experiments, we follow Song et al. (2023); Wang et al. (2021) and use the pre-trained WideResNet-28 (Zagoruyko & Komodakis, 2016) model from the RobustBench benchmark (Croce et al., 2021). The latent representation have a dimension of 512.
  • DenseNet: For FMoW-Time, we follow the backbone choice of Koh et al. (2021); Yao et al. (2022) and use DenseNet121 (Huang et al., 2017) for the land use classification task. Weights for three random trainings seeds are provided by Yao et al. (2022). We use the checkpoints for plain empirical risk minimization. The latent representation dimension is 1024.
  • ResNet: For EVIS, we follow Zhou et al. (2022) and use their ResNet-18 (He et al., 2016) model with a representation dimension of 512 and train on three random seeds. For ImageNet, we follow Song et al. (2023) and employ the standard pre-trained ResNet-50 model from RobustBench (Croce et al., 2021). Latent representations are of dimension 2048.
C.2 Baselines
  • Source Model: the un-adapted original model.
  • BatchNorm (BN) Adaptation (Schneider et al., 2020; Nado et al., 2020): aims to adapt the source model to distributions shift by collecting normalization statistics (mean and variance) of the test data.
  • Test Entropy Minimization (TENT) (Wang et al., 2021): goes one step further and optimizes the BN transformation parameters (scale and shift) by minimizing entropy on test predictions.
  • Continual Test-Time Adaptation (CoTTA) (Wang et al., 2022): takes a different approach by optimizing all model parameters with an entropy objective on augmentation averaged predictions and combines it with stochastic weight restore to prevent catastrophic forgetting.
  • Source HypOthesis Transfer (SHOT) (Liang et al., 2020) adapts the feature extractor via an information maximization loss in order to align the representations with the source classifier.
  • Sharpness-Aware Reliable Entropy Minimization (SAR) (Niu et al., 2023) filters out samples with large gradients based on their entropy values and encourages convergence to a flat minimum.
  • Laplacian Adjusted Maximum likelihood Estimation (LAME) (Boudiaf et al., 2022) regularizes the likelihood of the source model with a Laplacian correction term that encourages neighbouring representations to be assigned to the same class.
  • Test-Time Template Adjuster (T3A) (Iwasawa & Matsuo, 2021) computes new class prototypes by a running average of low entropy representations.
C.3 Implementation Details and Hyperparameters
By the nature of test-time adaptation, choosing hyperparameters is tricky (Zhao et al., 2023b) since one cannot assume access to a validation set of the test distribution in practise. To ensure we report the optimal performance on new or barely used datasets (Yearbook, EVIS, FMoW, CIFAR-10.1 and ImageNetV2), we perform a grid search over hyperparameters as suggested in the original papers. We perform separate grid searches for the uniform label distribution and online imbalanced label distribution setting. Reported performance correspond to the best setting. If the baselines were studied in the gradual CIFAR-10-C setting by Wang et al. (2022), we use their hyperparameter setup; otherwise, we conduct a grid search as described earlier. Unless there is a built-in reset (SAR) or convergence criteria (LAME) all methods run without reset and one optimization step is performed. We use the same batch sizes for all baselines. For Yearbook we comprise all samples of a year in one batch resulting in a batch size of 2048. To create online class imbalance, we reduce the batch size to 64. We use a batch size of 100 for EVIS, CIFAR.10.1 and CIFAR-10-C and 64 for FMoW-Time and ImageNetV2.